# Theory of two-component superfluidity of microcavity polaritons

### v1

@@ II Effective Hamiltonian of upper and lower polaritons
$=\sum_{\mathbf{P}}\varepsilon_{\text{LP}}(\mathbf{P})\hat{l}_{\mathbf{P}}^{\dagger}\hat{l}_{\mathbf{P}}+\sum_{\mathbf{P}}\varepsilon_{\text{UP}}(\mathbf{P})\hat{u}_{\mathbf{P}}^{\dagger}\hat{u}_{\mathbf{P}}$
$\quad+\frac{1}{2A}\sum_{\begin{subarray}{c}\mathbf{P},\mathbf{P}^{\prime},\mathbf{q}\end{subarray}}U_{q}\cdot\frac{1}{4}\Big{[}\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}$
$\qquad+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}$
$\qquad+\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}\Big{]}.$
@@ III Spectrum of Collective Excitations
We expect that at $T=0$ K, almost all upper and lower polaritons condense into a BEC of upper and lower polaritons. This allows us to assume the formation of a binary mixture of BECs. We describe this two-component weakly interacting Bose gas of polaritons following the procedure, described in Refs. SuperfluidDipolar ; HighTempSuperOleg ; Pindzola . We then examine this mixture within the Bogoliubov approximation Abrikosov . The collective excitation spectrum for a two-component weakly interacting Bose gas of polaritons is given by
$\Omega_{j,\mathbf{P}}=\sqrt{\frac{\omega_{1,\mathbf{P}}^{2}+\omega_{2,\mathbf{P}}^{2}+(-1)^{j-1}\sqrt{\left(\omega_{1,\mathbf{P}}^{2}-\omega_{2,\mathbf{P}}^{2}\right)^{2}+\left(4G_{12}\right)^{2}\varepsilon_{1,\mathbf{P}}\varepsilon_{2,\mathbf{P}}}}{2}},$ (17)
where $\omega_{j,\mathbf{P}}=\sqrt{\varepsilon_{j,\mathbf{P}}^{2}+2G_{jj}\varepsilon_{j,\mathbf{P}}}$, with $j=1$ for lower polaritons (LP) and $j=2$ for upper polaritons (UP), and their spectra are defined in Eq. (60). Self-interactions are dictated as follows,
$G_{jj}=U_{q}n_{j,0}\ ,$ (18)
[…]
The cross-interaction term is defined as follows,
$G_{12}=U_{q}\sqrt{n_{\text{LP},0}n_{\text{UP},0}}.$ (19)
We expand the spectrum of collective excitations $\Omega_{j,\textbf{P}}$ up to first order with respect to the momentum P and get two modes of the collective excitations $\Omega_{j,\textbf{P}}=c_{j}\textbf{P}$.
$c_{j}=\sqrt{\frac{G_{\text{11}}}{2m_{\text{LP}}}+\frac{G_{\text{22}}}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{G_{\text{11}}}{2m_{\text{LP}}}-\frac{G_{\text{22}}}{2m_{\text{UP}}}\right)^{2}+\frac{G_{\text{12}}^{2}}{m_{\text{LP}}m_{\text{UP}}}}}\ ,$ (20)
where $j=1,2$, and $G_{11},G_{22}$ and $G_{12}$ are polariton-polariton interactions defined in Eq. (18) and Eq. (19). The upper and lower polariton masses $m_{\text{UP}}$ and $m_{\text{LP}}$ are defined in Eq. (71) and (70). If we consider the finely tuned and experimentally plausible case where the densities of the polaritons are equal, $n_{\text{LP}}=n_{\text{UP}}=n/2$, we can obtain the sound velocity as
$c_{j}=\sqrt{\frac{U_{q}n}{2}\left(\frac{1}{2m_{\text{LP}}}+\frac{1}{2m_{\text{UP}}}+(-1)^{j-1}\sqrt{\left(\frac{1}{2m_{\text{LP}}}-\frac{1}{2m_{\text{UP}}}\right)^{2}+\frac{1}{m_{\text{LP}}m_{\text{UP}}}}\right)}\ ,$ (21)
It is clear to see that in Eq. (21) there is only one non-zero sound velocity.
$c=\sqrt{\frac{U_{q}n}{2}\left(\frac{1}{m_{\text{UP}}}+\frac{1}{m_{\text{LP}}}\right)}\ .$ (22)
To account for the differences in the concentrations of upper and lower polaritons, we incorporate the Hopfield coefficients $X_{P}$ and $C_{P}$, which represent the excitonic and photonic fractions, respectively. The total Hamiltonian is expressed as,
$\hat{H}_{\text{total}}=\hat{H}_{0}+\hat{H}_{\text{int}}.$ (23)
The non-interacting Hamiltonian is given by
$\hat{H}_{0}=\sum_{\mathbf{P}}\varepsilon_{\text{LP}}(\mathbf{P})\hat{l}_{\mathbf{P}}^{\dagger}\hat{l}_{\mathbf{P}}+\sum_{\mathbf{P}}\varepsilon_{\text{UP}}(\mathbf{P})\hat{u}_{\mathbf{P}}^{\dagger}\hat{u}_{\mathbf{P}},$ (24)
where the energy spectra for lower and upper polaritons are given by Eq. (66). Here, the photon and exciton dispersions are given by Eq. (62)-(63). The Hopfield coefficients $X_{P}$ and $C_{P}$ determine the excitonic and photonic fractions, respectively and we restate them in a convenient form below Griffin1995 ,
$|X_{P}|^{2}=\frac{1}{2}\left(1+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4(\hbar\Omega_{R})^{2}}}\right),\quad|C_{P}|^{2}=\frac{1}{2}\left(1-\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4(\hbar\Omega_{R})^{2}}}\right).$ (25)
@@
We define $\Delta_{0}=\frac{c}{n}\hbar\pi L_{c}^{-1}-\left(E_{\text{band}}-E_{\text{binding}}\right)$ as the detuning. The Hopfield coefficients can be varied widely by varying the detuning and the Rabi splitting as seen below in Fig. 1.
@@
Figure 1: The Hopfield coefficients as a function of the detuning $\Delta_{0}$ defined above for GaAs. We employed the definitions in Eq. (25). We kept the Rabi splitting constant at $\Omega_{R}=$ 15 meV.
It is worth noting that when the detuning is set to 0 meV, the photon and exciton fractions are necessarily fixed at 0.5 each. At this particular detuning level, the Rabi splitting has no effect on the photon and exciton fraction. At any level of non-zero detuning, the fractions are nearly 100% photonic or 100% excitonic. At each of these extremes the detuning is the more prominent parameter than Rabi splitting. However, at a detuning close to, but not quite even, the Rabi splitting has a more prominent effect. This is seen below for GaAs in Fig. 2.
Figure 2: The Hopfield coefficients as a function of the Rabi splitting $\Omega_{R}$ defined above for GaAs. We employed the definitions in Eq. (25). We kept the detuning constant at $\Delta_{0}=$ 10 meV.
When the Rabi splitting is increased, the exciton fraction goes down, and the photon fraction increases. This coupled with the change of the effective masses of the polaritons with the Rabi splitting illustrated in Fig. 20 govern the dynamics of our system, and we will come to see, lead to rich dynamics of the sound velocity and critical temperatures when parametrized by the detuning and the Rabi splitting.
To simplify the notation, we absorb the interaction strength $U_{q}$ into the coefficients $G_{\text{LP}},G_{\text{UP}},$ and $G_{\text{LP-UP}}$. To reflect the asymmetry in polariton populations, we define $n_{\text{LP},0}$ and $n_{\text{UP},0}$ as the total polariton concentrations of lower and upper polaritons, respectively,
$n_{\text{LP},0}=n_{0}|X_{P}|^{2},\quad n_{\text{UP},0}=n_{0}|C_{P}|^{2},$ (26)
where $n_{0}$ is the total polariton concentration. These redefinitions result in the following interaction strengths,
$G_{\text{LP}}=U_{q}n_{0}|X_{P}|^{2},\quad G_{\text{UP}}=U_{q}n_{0}|C_{P}|^{2},\quad G_{\text{LP-UP}}=U_{q}n_{0}|X_{P}||C_{P}|.$ (27)
The interaction Hamiltonian then becomes,
$\hat{H}_{\text{int}}=\frac{1}{2A}\sum_{\begin{subarray}{c}\mathbf{P},\mathbf{P}^{\prime},\mathbf{q}\end{subarray}}\Big{[}G_{\text{LP}}\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{l}_{\mathbf{P}^{\prime}}+G_{\text{UP}}\hat{u}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+G_{\text{LP-UP}}\left(\hat{l}_{\mathbf{P}+\mathbf{q}}^{\dagger}\hat{u}_{\mathbf{P}^{\prime}-\mathbf{q}}^{\dagger}\hat{l}_{\mathbf{P}}\hat{u}_{\mathbf{P}^{\prime}}+\text{.....}\right)\Big{]}.$.. ) ] . (28)
This modifies the interaction terms, leading to observable effects in the collective excitation spectrum and sound velocities, which we derive in the next section.
IV Tunable sound velocity
The collective excitation spectrum for the two-component polariton condensate is given by
$\varepsilon_{j,{P}}=\frac{1}{\sqrt{2}}\left\{\varepsilon_{\text{LP}}({P})^{2}+2G_{\text{LP}}\varepsilon_{\text{LP}}({P})+\varepsilon_{\text{UP}}({P})^{2}+2G_{\text{UP}}\varepsilon_{\text{UP}}({P})+(-1)^{j-1}\sqrt{\beta({P})}\right\}^{1/2}\ ,$ (29)
where
$\beta({P})=\left[\varepsilon_{\text{LP}}({P})^{2}+2G_{\text{LP}}\varepsilon_{\text{LP}}({P})-\varepsilon_{\text{UP}}({P})^{2}-2G_{\text{UP}}\varepsilon_{\text{UP}}({P})\right]^{2}+16G_{\text{LP-UP}}^{2}\varepsilon_{\text{LP}}({P})\varepsilon_{\text{UP}}({P}).$ (30)
We plot this collective excitation spectrum in Fig. 3.
Figure 3: The spectrum of collective excitations in GaAs as a function of momentum. The parameters used in the calculation are $\Omega_{R}=17$ meV, a bare photon of energy 1.6 eV, and a finely adjusted detuning quantity $\Delta_{0}=0$ meV. The collective excitation spectrum is defined in Eq.  (29).
A notable feature worth pointing out about Fig. 3 is the shape of the curves. The collective excitation spectrum for the j = 1 mode is linear in the low momentum regime. This is the excitation which corresponds to the Goldstone mode for superfluidity, and thus the non-zero sound velocity $c_{1}$ in Eq. (39). This mode arises as a result of the spontaneous breaking of the U(1) symmetry of the condensate. Since the condensate prefers a specific phase, small fluctuations in this phase propagate as low-energy excitations, leading to a gapless, sound-like dispersion. In contrast, the j = 2 mode represents a density fluctuation of the condensate.
We can simplify Eq. (29) to give
$\varepsilon_{j,{P}}=\sqrt{\frac{\omega_{\text{LP},{P}}^{2}+\omega_{\text{UP},{P}}^{2}+(-1)^{j-1}\sqrt{\beta({P})}}{2}},$ (31)
where,

### v2

@@ II Effective Hamiltonian of upper and lower polaritons
$=\sum_{P}\varepsilon_{\text{LP}}(P)\,\hat{l}_{P}^{\dagger}\hat{l}_{P}\;+\;\sum_{P}\varepsilon_{\text{UP}}(P)\,\hat{u}_{P}^{\dagger}\hat{u}_{P}$
$\quad+\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}U_{q}\Bigg\{\,g_{\ell\ell}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}$
$\qquad\qquad\qquad+\,g_{uu}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$
$\qquad\qquad\qquad+\,g_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}$
$\qquad\qquad\qquad+\,g_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$\qquad\qquad\qquad+\,\tilde{g}_{\ell u}^{PP^{\prime}q}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$\qquad\qquad\qquad+\,\tilde{g}_{u\ell}^{PP^{\prime}q}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}\Bigg\}.$
We emphasize that the interaction coefficients $g_{\ell\ell},g_{uu},g_{\ell u},g_{u\ell},\tilde{g}_{\ell u},\tilde{g}_{u\ell}$ defined above are in general momentum-dependent through the Hopfield fractions $X_{P}$ and $C_{P}$. Retaining the full momentum dependence ensures that inter- and intra-branch scattering are treated rigorously across all detunings and Rabi couplings. However, such generality comes at the cost of analytic tractability: closed-form expressions for the Bogoliubov spectrum, sound velocity, and critical temperature become intractable if one carries the momentum dependence throughout. For this reason, we adopt a controlled approximation appropriate for condensates forming near $P=0$. We thus evaluate the Hopfield coefficients at zero in-plane momentum, $X_{P}\to X_{0}$ and $C_{P}\to C_{0}$, and use these values to define effective interaction strengths. This approximation is physically justified because Bose-Einstein condensation occurs predominantly at zero momentum, where the single-particle energies are minimized and superfluid properties such as sound velocity and critical temperature are governed by low-energy, long-wavelength collective modes rather than high-momentum excitations. Additionally, at small $P$, the Hopfield coefficients vary slowly, so $X_{P}\approx X_{0}$ and $C_{P}\approx C_{0}$ remain accurate.
@@ III Spectrum of collective excitations
In analyzing the excitation spectrum of the coexisting LP-UP condensate, we may begin in the long-wavelength limit ($P\to 0$) where the dispersions can be linearized and the effective polariton masses treated as constants. This is the regime directly relevant for superfluid properties such as the sound velocity and the critical temperature. Both quantities are governed by the low-energy, phonon-like Bogoliubov modes rather than by high-momentum single-particle excitations. Moreover, at long wavelengths the momentum dependence of the Hopfield coefficients becomes negligible, allowing the interaction Hamiltonian to be expressed in terms of constant effective couplings AGD2012 . This approximation preserves the essential collective physics while avoiding artifacts introduced by momentum-dependent mass renormalization, and is standard practice when applying the Landau criterion or deriving analytic estimates for $c_{s}$ and $T_{c}$.
@@
We define $\Delta_{0}=\frac{c}{n}\hbar\pi L_{c}^{-1}-\left(E_{\text{band}}-E_{\text{binding}}\right)$ as the detuning. At $P=0$ the Hopfield coefficients become momentum independent,
$X_{P}\to X_{0},\qquad C_{P}\to C_{0},\quad X_{0}^{2}=\tfrac{1}{2}\!\left(1+\frac{\Delta_{0}}{\sqrt{\Delta_{0}^{2}+4(\hbar\Omega_{R})^{2}}}\right),\ C_{0}^{2}=1-X_{0}^{2},$ (23)
Crucially, the Hopfield coefficients retain their full detuning and Rabi-splitting dependence. This allows us to explore off-resonance regimes while maintaining analytic tractability. The resulting expressions for collective modes, sound velocity, and critical temperature are exact within the long-wavelength limit and provide physically faithful benchmarks for comparison with experiments and future numerical treatments that incorporate full momentum dependence. The Hopfield coefficients at $P=0$ can be varied widely by varying the detuning and the Rabi splitting as seen below in Figure 1.
@@
Figure 1: The Hopfield coefficients as a function of the detuning $\Delta_{0}$ defined above for WSe2. We employed the definitions in Eq. (23). We kept the Rabi splitting constant at $\Omega_{R}=$ 8 meV.
For the contact interaction we take $U_{q}\to U_{0}$ using Eq. (12) for GaAs/AlGaAs quantum wells, and the experimentally measured value from polariton blueshift measurements for monolayers of WSe2  Kravtsov2020 ; flatten2016 . We then define the constant effective couplings at $P=0$
$g_{\ell\ell}\equiv U_{0}|X_{0}|^{4},\qquad g_{uu}\equiv U_{0}|C_{0}|^{4},\qquad g_{\ell u}\equiv U_{0}|X_{0}|^{2}|C_{0}|^{2},$ (24)
the interaction Hamiltonian in the polariton basis reduces to
$\hat{H}_{\text{int}}^{(P=0)}=\frac{1}{2A}\sum_{\begin{subarray}{c}P,P^{\prime},q\end{subarray}}\Big[$ $g_{\ell\ell}\,\hat{l}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{l}_{P^{\prime}}+g_{uu}\,\hat{u}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{u}_{P^{\prime}}$ (25)
$+g_{\ell u}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}+g_{\ell u}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}$
$+g_{\ell u}\,\hat{l}_{P+q}^{\dagger}\hat{u}_{P^{\prime}-q}^{\dagger}\hat{u}_{P}\hat{l}_{P^{\prime}}+g_{\ell u}\,\hat{u}_{P+q}^{\dagger}\hat{l}_{P^{\prime}-q}^{\dagger}\hat{l}_{P}\hat{u}_{P^{\prime}}\Big].$
The non-interacting contribution to the total Hamiltonian is given by Eq. (11), where the single-particle dispersions $\varepsilon_{\mathrm{LP/UP}}$ are defined in Appendix B, Eq. (65). Collecting all terms, we obtain
$\boxed{\hat{H}_{\text{total}}^{(P=0)}=\hat{H}_{0}+\hat{H}_{\text{int}}^{(P=0)}},$ (26)
we obtain the long-wavelength effective Hamiltonian we will use to derive the collective excitation spectrum.
At $T=0$ K we assume that both lower and upper polaritons macroscopically occupy the zero-momentum state, forming a coexisting condensate. Expanding the total Hamiltonian in fluctuations around the condensates and applying the Bogoliubov approximation, we obtain the quadratic Hamiltonian governing the collective excitations. Diagonalization yields two excitation branches with dispersion
$\Omega_{j,P}=\sqrt{\frac{\omega_{\text{LP},P}^{2}+\omega_{\text{UP},P}^{2}+(-1)^{j-1}\sqrt{\left(\omega_{\text{LP},P}^{2}-\omega_{\text{UP},P}^{2}\right)^{2}+\left(4G_{\text{LP-UP}}\right)^{2}\varepsilon_{\text{LP}}(P)\varepsilon_{\text{UP}}(P)}}{2}}\ ,$ (27)
where $j=1,2$ and